OpenStudy (anonymous):

11x-77/9 divided by y+1/y+3 11x-77/9 divided by y+1/y+3 @Mathematics

OpenStudy (anonymous):

dont i flip the second one?

OpenStudy (anonymous):

\[\frac{11x-77}{9}\div \frac{y+1}{y+3}=\frac{11x-77}{9}\times \frac{y+3}{y+1}\]

OpenStudy (anonymous):

yes flip the second one

OpenStudy (anonymous):

oh wait i wrote the wrong thing im sorry

OpenStudy (anonymous):